OpenStudy (j.f):

What is the probability that it will rain on a given day if the probability that it will not rain is 1/3? If there are p ways the first way can be made, and q ways a second can be made, then pxq 1/3x1/7 (7 days in a week) 1/21?

Directrix (directrix):

Probabiility that it will rain + probability it will NOT rain = 1 Prob of rain = 1 - 1/3 = 2/3

OpenStudy (anonymous):

Are you trying to find the probability that it will rain in a week or a day?

Directrix (directrix):

This is what I am doing and thought that the question meant: What is the probability that it will rain on a given day.

Directrix (directrix):

Have you studied binomial probability yet?

OpenStudy (j.f):

Yes, polynomials and all that good stuff. I figured the meant 7 days with "on a given day"

OpenStudy (j.f):

But you are right